OpenStudy (anonymous):

Multiply these binomials. (x - 9)(x + 9)

OpenStudy (callisto):

(x - 9)(x + 9) = x(x + 9) -9(x + 9) = x^2 +9x - 9x -81 =x^2 -81

OpenStudy (hoblos):

you can use the formula of difference of two squares directly.. (a-b)(a+b) = a^2 - b^2